# Weakly supervised anomaly detection with event-level variables

### v1

@@ Abstract
We introduce a new topology for weakly supervised anomaly detection searches, di-object plus X. In this topology, one looks for a resonance decaying to two standard model particles produced in association with other anomalous event activity (X). This additional activity is used for classification. We demonstrate how anomaly detection techniques which have been developed for di-jet searches focusing on jet substructure anomalies can be applied to event-level anomaly detection in this topology. To robustly capture event-level features of multi-particle kinematics, we employ new physically motivated variables derived from the geometric structure of a collision’s phase space manifold. As a proof of concept, we explore the application of this approach to several benchmark signals in the di-$\tau$ plus X final state. We demonstrate that our anomaly detection approach can reach discovery-level significances for signals that would be missed in a conventional bump-hunt approach.
@@ I Introduction
The majority of proposed anomaly detection algorithms can generally be split into two separate categories: unsupervised and weakly supervised algorithms. Unsupervised algorithms, first proposed in Refs. [11, 12], are trained solely on background events and are then used to define an anomaly score which quantifies how similar an input event is to typical background events. Weakly supervised methods build on the classification without labels paradigm [13] to learn to distinguish between signal and background. This is accomplished by training a classifier to distinguish events in the signal region (SR), which may contain some small fraction of signal, from a carefully constructed sample of background-only events [14, 15]. Weakly supervised methods have the advantage that they are asymptotically optimal: the optimal weakly supervised classifier will learn the likelihood ratio, which is the ideal signal versus background classifier.
@@
Because weakly supervised algorithms need a high-quality estimate of the background in a high-dimensional space, their applications have mostly focused on resonant signals, so the background composition in a signal mass window can be learned through an interpolation from sideband regions. Several approaches have been proposed to achieve the necessary high-quality background estimate [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. The majority of these first anomaly detection approaches have focused on heavy resonances decaying to two boosted large radius jets, where well-studied jet substructure observables, with minimal correlations with the resonance mass, can be used as the classification features. There is also a large space of signal models which could produce many phenomenological varieties of jets, making this topology ideal for anomaly detection. However, considerably less effort has been put into exploring other signal topologies which could work with a weakly supervised resonant anomaly detection approach.
Several works have studied the application of anomaly detection methods to topologies beyond di-jet resonances. Refs. [24, 25] studied the application of weak supervision to topologies with high missing energy and a boosted jet. Ref. [26] studied the application of CATHODE to supersymmetry scenarios producing pairs of gluinos. Recently, Ref. [10] applied the CATHODE method to CMS open data to demonstrate the identification of $\Upsilon$ particles decaying to muon pairs. This study used only features related to the dimuon resonance for classification rather than focusing on the other event activity we exploit here. A previous ATLAS search [6] applied an event-level unsupervised autoencoder-based anomaly detection algorithm to search for resonances for several di-object combinations, where at least one of the objects was required to be a jet.
@@
In this paper, we explore the application of weakly supervised anomaly detection in a new di-object + X topology using the CATHODE methodology [18] with event-level features. We consider signals which produce a two-body resonance in addition to other particles (X). Such a signature can easily arise if the resonance is produced through an extended decay chain or from associated production with additional particles. These production mechanisms can lead to striking experimental signatures, with many objects in the event, potentially with rich kinematic correlations, and/or high missing energy. These distinct signatures can be leveraged to significantly reduce backgrounds; however, given the large phenomenological space of possibilities it is infeasible to design analysis strategies for every possible scenario. These two factors make for an ideal application of anomaly detection methods. This work develops a general framework for weakly supervised searches using event-level classification. The usage of weakly supervised, rather than unsupervised, anomaly detection may also offer larger discovery potential for anomalies due to improved classification performance [27].
@@
As a first exploration of this general topology, this work focuses specifically on the scenario where the resonance decays into two $\tau$ leptons. Searches featuring $\tau$ leptons are strongly motivated by various BSM theories, which often preferentially couple to third-generation fermions, but are less explored at the LHC due to their more complex experimental signatures. Examples of searches include $Z^{{}^{\prime}}$ bosons [28], leptoquarks [29, 30, 31], and additional Higgs bosons [32, 33]. These new particles are motivated by the observed anomalies in $B$-hadron decays to $D^{(\star)}\tau\nu$, which show enhanced rates compared to the corresponding decays involving lighter leptons [34, 35, 36, 37, 38, 39, 40, 41, 42]. If these anomalies are caused by heavy BSM particles, it is likely they also influence the di-$\tau$ production at high invariant masses.
@@
Though we focus on the di-$\tau$ final state in this work, the methodology we develop is entirely general and is applicable to other resonance + X searches, such as di-muon, di-electron, di-photon, or di-jet resonance + X. Although these other final states may have somewhat better existing experimental coverage, there is almost certainly still unexplored ‘X’ signatures which could be covered by our approach. These other final states may also offer the advantage of larger samples available for the training of the background model of the CATHODE method which can improve performance.
@@
The simulated data used in this paper can be found under Refs. [43, 44] and the analysis code is available on GitHub [45].
@@ II Setup
In this study, we focus on the final state where both $\tau$ leptons decay hadronically, which is the decay mode with the largest branching ratio. Unlike di-jet searches where the QCD multijet background is dominant, di-$\tau$ searches receive considerable backgrounds from multiple physics processes. The dominant backgrounds in the all hadronic final state are Drell–Yan and $\mathrm{t}\bar{\mathrm{t}}$ QCD and W+jets also contribute in this channel, with the fraction dependent on the purity of the $\tau$ identification criteria being employed and pre-selection requirements. As this is an exploratory work, and these effects are not modeled well with simulation, we focus here on the Drell–Yan and $\mathrm{t}\bar{\mathrm{t}}$ backgrounds only. The methodology we employ should generally account for any background which varies smoothly in di-$\tau$ mass. Due to the multiple decay modes of the $\tau$ and limited reconstruction efficiencies, di-$\tau$ searches feature significantly smaller event samples than the di-jet searches of prior weakly supervised searches. This presents an interesting challenge for the application of weakly supervised methods which require a large sample of background events in order to properly learn the probability distribution of the background events.
We apply the CATHODE method for resonant anomaly detection [18]. In CATHODE  one defines a resonant variable, in our case the visible mass of the di-$\tau$ candidate, $m_{\mathrm{vis}}$, and a set of features to be used for classification, $\vec{x}$. Signals are assumed to manifest in a localized region of the resonant variable, allowing the probability distribution of the background to be learned from neighboring sideband regions. The method proceeds in two steps. First, a generative model is trained on the sideband regions to learn the probability density of $\vec{x}$ conditioned on $m_{\mathrm{vis}}$ for the background, $P_{b}(\vec{x}|m_{\mathrm{vis}})$. The trained generative model is then interpolated into the signal region and used to generate synthetic background events. A weakly supervised classifier is then trained to distinguish between the real data events $p_{\mathrm{data}}(\vec{x})$ in the signal region and the synthetic background events $p_{\mathrm{bkg}}(\vec{x})$. If there is a true signal in the signal region abundant in sufficient quantity, then the classifier will learn to identify its events, which will look anomalous with respect to the interpolated background.
@@
For these reasons, we explore the usage of the recently introduced [49, 50] phase space variables to encode the relevant kinematic features of the event in a compact representation. In Ref. [50] a covariant description of the phase space of collider events producing a chosen number ($N$) of massless final-state particles was presented. The manifold of this $N$-body phase space, $\Pi_{N}$, was shown to be isomorphic to the product of an ($N-1$)-simplex ($\Delta_{N-1}$) and a ($32N-$ 3)-hypersphere ($S_{2N-3}$), i.e. $\Pi_{N}\cong\Delta_{N-1}\times S_{2N-3}$. A set of explicit and global coordinates of the phase space have been introduced: $(N)$ $\rho$ coordinates on the simplex $\Delta_{N-1}$ and $(N-1)$ complex $\xi$ coordinates on the hypersphere $S_{2N-3}$. Ref. [50] used this representation to develop a distance metric between events produced at a hadron collider using these global phase space coordinates. In this work, we do not use this distance metric, but rather directly use the phase space coordinates as a way to represent the kinematics of all the final state objects in the event. We exploit the simplex coordinates only, since we found that the normalizing flows used for the background model, described in Section IV, have difficulties with the non-trivial topology of the hypersphere coordinates.
@@ III Simulated Samples
For the SM background samples we used the SM UFO [56] with CKM mixing, and with the four-flavor scheme. We simulated extra jets (ISR/FSR) up to 2 jets, i.e. 0, 1, or 2 jets besides the hard scatter. Matrix element and parton shower matching was performed using the MLM scheme [57] with a matching scale of 30 GeV. For the $t\bar{t}$ sample, the top quarks were decayed in MadSpin [58] to W bosons that decay to $\tau$ leptons. In order to enhance the generation efficiency, we applied a generator level cut on the di-$\tau$ invariant mass of 120 GeV. The cross section of the Drell–Yan sample is $17.37$ pb and the cross section of the $t\bar{t}$ sample is $13.49$ pb. Sufficient events were generated to have an equivalent luminosity of 138 fb-1, to match that of the proton-proton collisions data taken during Run-2 (2016-18) of the LHC.
We explored a set of several BSM models as benchmarks to evaluate the performance of our approach. The details of the generation are summarized in Table 1, where the number of events is after the generator cuts are applied. Some example Feynman diagrams of the signal processes are shown in Fig. 2. For simplicity, we set the masses of the di-$\tau$ resonance in all models to 250 GeV.
In the 2HDM model [59, 60, 61] we use the heavy Higgs particle, denoted with $\Phi$, and let it decay to a $\tau$ lepton pair. In the eVLQ model [62, 63], we let the vector-like quark, $T^{\prime}$ decay to a SM top and a BSM scalar $S^{0}$. One of the $S^{0}$ then decays to a $\tau$ lepton pair while the other decays inclusively according to branching ratios similar to a heavy Higgs boson. In the NMSSM model [64, 65, 66, 67, 68] we simulate cascade processes, where the heavy resonance X decays to a Higgs boson (H) and an unknown particle Y. Y further decays to the $\tau$ pair and the H decays invisibly, leading to missing transverse momentum. We explored non-resonant signals such as vector-like anti-leptons [69], leptoquarks [70] and heavy neutral leptons [71]. We did not achieve significant improvements with our method on these signals, as expected since these models are not well localized to the signal region, making them difficult to identify with the CATHODE approach. We therefore did not study them further.
@@
Figure 1 shows the reconstructed invariant visible mass distribution of the di-$\tau$ system. While the generator level mass cut is at 120 GeV, because some of the di-$\tau$ momentum is lost to the neutrinos, the reconstructed visible mass peaks at 100 GeV. Histograms are normalized to unit area.
Figure 1: Invariant visible mass distribution of the di-$\tau$ system.
Figure 2: Feynman diagrams of the $t\bar{t}\phi$ (left), $T^{\prime}\bar{T^{\prime}}$ (middle), and $X\to YH$ (right) signal models. The ta symbol denotes the $\tau$ leptons, and $\sim$ is used to denote the anti-particles. The $S^{0}$ decays inclusively according to its branching ratios.The $H$ boson in the $X\to YH$ model decays invisibly.
UFO Name Extra jets Parameters #Events
SM Drell–Yan $\tau\bar{\tau}$ up to 2 N/A 3M
SM $t\bar{t}\to W^{-}W^{+}b\bar{b}\to\tau\bar{\nu}+\bar{\tau}\nu$ up to 2 N/A 2M
2HDM $t\bar{t}\Phi\to\tau\bar{\tau}$ 0 $m_{\Phi}=250$ GeV 100k
eVLQ $T^{\prime}\bar{T}^{\prime}\to tS^{0}\bar{t}S^{0}\to\tau\bar{\tau}+\text{incl}$ 0 $m_{T^{\prime}}=1000$ GeV $m_{S}=250$ GeV 100k
NMSSM $X\to YH\to\tau\bar{\tau}+\text{inv}$ 0 $m_{X}=800$ GeV $m_{H}=400$ GeV $m_{Y}=250$ GeV 100k
Table 1: List of samples used, and the details of the production.

### v4

@@ Abstract
We introduce a new topology for weakly supervised anomaly detection searches, di-object plus X. In this topology, one looks for a resonance decaying to two standard model particles produced in association with other anomalous event activity (X). This additional activity is used for classification. We demonstrate how anomaly detection techniques which have been developed for di-jet searches focusing on jet substructure anomalies can be applied to event-level anomaly detection in this topology. To robustly capture event-level features of multi-particle kinematics, we employ new physically motivated variables derived from the geometric structure of a collision’s phase space manifold. As a proof of concept, we explore the application of this approach to several benchmark signals in the di-$\tau$ and di-$\mu$ plus X final states. We demonstrate that our anomaly detection approach can reach discovery-level significances for signals that would be missed in a conventional bump-hunt approach.
@@ I Introduction
The majority of proposed anomaly detection algorithms can generally be split into two separate categories: unsupervised and weakly supervised algorithms. Unsupervised algorithms, first proposed in Refs. [11, 12], are trained solely on background events and are then used to define an anomaly score which quantifies how similar an input event is to typical background events. Weakly supervised methods build on the classification without labels paradigm [13] to learn to distinguish between signal and background. This is accomplished by training a classifier to distinguish events in the signal region, which may contain some small fraction of signal, from a carefully constructed sample of background-only events [14, 15]. Weakly supervised methods have the advantage that they are asymptotically optimal: the optimal weakly supervised classifier will learn the likelihood ratio, which is the ideal signal versus background classifier [13].
@@
The training procedure of weakly supervised anomaly detection methods requires a high quality estimate of the background in the high-dimensional feature space, allowing the classifier to identify the anomalous data events in the signal region which are distinct from background. Defects in the background estimate will cause the classifier to learn the difference between the true background events in the signal region and the estimate, rather than identifying the signal. This requirement means that simulations, which are known to have minor mismodelings, are generally not suitable for such a background estimate and data-driven methods must be used. Most weakly supervised anomaly detection methods have therefore focused on resonant signals, so the background composition in a signal mass window can be learned through an interpolation from sideband regions. Several approaches have been proposed to achieve the necessary data-driven high-quality background estimate [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Many of these first anomaly detection approaches have focused on heavy resonances decaying to two boosted large radius jets, where well-studied jet substructure observables, with minimal correlations with the resonance mass, can be used as the classification features. There is also a large space of signal models which could produce many phenomenological varieties of jets, making this topology ideal for anomaly detection. The weakly supervised searches from at the LHC [4, 7, 8] have also targeted this dijet topology. Now that these first searches have been performed, it is timely to explore other signal topologies which are amenable to a weakly supervised resonant anomaly detection approach.
@@
In this paper, we explore the application of weakly supervised anomaly detection in a new di-object + X topology using the CATHODE methodology [18] with event-level features. The CATHODE method follows the general strategy described earlier and leverages a generative model trained on sideband data to characterize background distributions, subsequently interpolating these learned representations into the signal region to produce realistic background samples for weakly supervised anomaly detection. We consider signals which produce a two-body resonance in addition to other particles (X). Such a signature can easily arise if the resonance is produced through an extended decay chain or from associated production with additional particles. These production mechanisms can lead to striking experimental signatures, with many objects in the event, potentially with rich kinematic correlations, and/or high missing energy. These distinct signatures can be leveraged to significantly reduce backgrounds; however, given the large phenomenological space of possibilities it is infeasible to design analysis strategies for every possible scenario. These two factors make for an ideal application of anomaly detection methods. This work develops a general framework for weakly supervised searches in this topology using event-level classification. The usage of weakly supervised, rather than unsupervised, anomaly detection may also offer larger discovery potential for anomalies due to improved classification performance [24].
@@
As an exploratory study, we examine the specific case where the resonance decays into charged lepton pairs ($\mu$ or $\tau$). Since we expect similar behavior between electrons and muons, we focus our analysis on $\mu$ pairs instead of electrons. Searches featuring $\tau$ leptons are strongly motivated by various BSM theories, which often preferentially couple to third-generation fermions, but are less explored at the LHC due to their more complex experimental signatures. Examples of searches include $Z^{{}^{\prime}}$ bosons [25], leptoquarks [26, 27, 28], and additional Higgs bosons [29, 30]. These new particles are motivated by the observed anomalies in $B$-hadron decays to $D^{(\star)}\tau\nu$, which show enhanced rates compared to the corresponding decays involving lighter leptons [31, 32, 33, 34, 35, 36, 37, 38, 39]. If these anomalies are caused by heavy BSM particles, it is likely they also influence the di-$\tau$ production at high invariant masses. Searches for $\mu$ lepton resonances are also strongly motivated by various BSM theories. While there are a plethora of experimental searches looking for BSM resonances decaying to second generation fermions [40, 41, 42, 43, 44, 45, 46], it is likely there are still many unexplored exotic signatures.
@@
Though we focus on the di-$\tau$ and di-$\mu$ final states in this work, the methodology we develop is entirely general and is applicable to other resonance + X searches, such as di-electron, di-photon, or di-jet resonance + X. Although these other final states may have somewhat better existing experimental coverage, there is almost certainly still unexplored ‘X’ signatures which could be covered by our approach. These other final states may also offer the advantage of larger samples available for the training of the background model of the CATHODE method which can improve performance.
Several other works have studied the application of anomaly detection methods to topologies beyond di-jet resonances. In Refs. [47, 48] weak supervision was applied to topologies with high missing energy and a boosted jet while in Refs. [49, 50] the application of CATHODE to supersymmetry scenarios producing pairs of gluinos was examined. Recently, Ref. [10] applied the CATHODE method to CMS open data to demonstrate the identification of $\Upsilon$ particles decaying to muon pairs. This study used only features related to the dimuon resonance for classification rather than focusing on the other event activity we exploit here A previous ATLAS search [6] applied an event-level unsupervised autoencoder-based anomaly detection algorithm to search for resonances for several di-object combinations, where at least one of the objects was required to be a jet. These analyses used specific characteristics of these signal topologies to design the data-driven background estimate necessary for weak supervision, making generalization to other topologies unclear. The resonance + X topology considered in this paper complements these prior works by offering a framework to apply weakly supervised anomaly detection to searches in many different final states.
@@
The simulated data used in this paper can be found under Refs. [51, 52, 53] and the analysis code is available on GitHub [54].
@@ II Setup
II.1 Final States
Unlike di-jet searches where the QCD multijet background is dominant, di-$\tau$ and di-$\mu$ searches receive considerable backgrounds from multiple physics processes. In the di-$\tau$ study, we focus on the final state where both $\tau$ leptons decay hadronically, which is the decay mode with the largest branching ratio. The dominant backgrounds in the all hadronic di-$\tau$ and the di-$\mu$ final states are Drell–Yan and $\mathrm{t}\bar{\mathrm{t}}$ . QCD and W+jets also contribute in this channel, with the fraction dependent on the purity of the $\tau$ and $\mu$ identification criteria being employed and pre-selection requirements. As this is an exploratory work, and these effects are not modeled well with simulation, we focus here on the Drell–Yan and $\mathrm{t}\bar{\mathrm{t}}$ backgrounds only. The methodology we employ should generally account for any background which varies smoothly in di-$\tau$ or di-$\mu$ mass. Due to the multiple decay modes of the $\tau$ and limited reconstruction efficiencies, di-$\tau$ searches feature significantly smaller event samples than the di-jet or di-$\mu$ searches of prior weakly supervised searches. This presents an interesting challenge for the application of weakly supervised methods which require a large sample of background events in order to properly learn the probability distribution of the background events.
II.2 The CATHODE Method
We apply the CATHODE method for resonant anomaly detection [18]. In CATHODE, one defines a resonant variable, in our case the visible mass of the di-$\tau$ candidate ($m_{\mathrm{vis}}$) or invariant mass of the di-$\mu$ candidate ($m_{\mu\mu}$), and a set of features to be used for classification, $\vec{x}$. Signals are assumed to manifest in a localized region of the resonant variable and backgrounds are assumed to vary smoothly as function of the resonant variable. These assumptions allow the probability distribution of the background to be learned from neighboring sideband regions. The method proceeds in two steps. First, a generative model is trained on the sideband regions to learn the probability density of $\vec{x}$ conditioned on $m$ for the background, $P_{b}(\vec{x}|m)$. The trained model is then interpolated into the signal region and used to generate synthetic background events. A weakly supervised classifier is then trained to distinguish between the real data events $p_{\mathrm{data}}(\vec{x})$ in the signal region and the synthetic background events $p_{\mathrm{bkg}}(\vec{x})$. If there is a true signal in the signal region abundant in sufficient quantity, then the classifier will learn to identify its events, which will look anomalous with respect to the interpolated background.
@@ II.2 The CATHODE Method
For these reasons, we explore the usage of the recently introduced [58, 59] phase space variables to encode the relevant kinematic features of the event in a compact representation.
II.3 Phase Space Variables
In Ref. [58, 59] a covariant description of the phase space of collider events involving a chosen number ($N$) of massless final-state particles was presented. The manifold of this $N$-body phase space, $\Pi_{N}$, was shown to be isomorphic to the product of an (N1N-1)-simplex ($\Delta_{N-1}$) and a (2N32N-3)-hypersphere ($S_{2N-3}$), i.e. $\Pi_{N}\cong\Delta_{N-1}\times S_{2N-3}$. A set of explicit global coordinates on the phase space manifold were introduced: (N1)(N-1) coordinates $\rho$ on the simplex $\Delta_{N-1}$ and (2N3)(2N-3) angular coordinates $\xi,\chi$ on the hypersphere $S_{2N-3}$. For $N=2$, the simplex (hypersphere) coordinates lie on a line (circle), while for $N=3$ they span a triangle (sphere). Ref. [59] used this representation to develop a metric between events produced at a hadron collider in terms of their phase space coordinates and demonstrated that the distance between events could be used effectively for event classification. Although the phase space coordinates are simply functions of the final-state particles’ four-momenta, they are the natural variables for distinguishing different physics processes since they define the space on which $S$-matrix elements live. In this work, we do not use the phase space metric itself, but rather directly use the phase space coordinates as an efficient representation of the kinematics of all the final state objects in the event. We exploit the simplex coordinates only, since we found that they significantly enhance sensitivity while the generative model used for the background model, described in Section IV, has difficulties with the non-trivial topology of the hypersphere coordinates detailed in  A.2.
At a hadron collider, in the final state’s center of mass the simplex coordinates take the simple form
$\rho_{i}=\frac{1}{Q}(E_{i}\pm p_{z,i})$ (1)
where $Q$ is the total energy of the event in the center of mass frame, the index $i=1,\dots,N$ runs over the final-state particles, and the choice of sign is a matter of convention. The constraint $\sum_{i}\rho_{i}=1$ implies that the $\rho_{i}$ are coordinates on an N1N-1 dimensional simplex. These simplex variables can be expressed in terms of rapidity $y$ and transverse momentum $p_{T}$ as
$\rho_{i}=\frac{p_{T,i}}{Q}e^{\pm y_{i}}$ (2)
whereas the azimuthal angles of final-state particles are contained entirely in the sphere coordinates; we refer to Ref. [59] for further details. The simplex coordinates are a useful compact combination of $p_{T}$ and $y$ insofar as the simplex contribution to the phase space distance between events is just the Euclidean distance between the simplex coordinates. Indeed, we find that the simplex coordinates enhance sensitivity compared to a feature set of the same dimension using $p_{T}$; see Appendix B for a detailed study.
@@ III Simulated Samples
For the SM background samples we used the SM UFO [65] with CKM mixing, and with the four-flavor scheme. We simulated extra jets (ISR/FSR) up to 2 jets, i.e. 0, 1, or 2 jets besides the hard scatter. Matrix element and parton shower matching was performed using the MLM scheme [66] with a matching scale of 30 GeV. For the $t\bar{t}$ sample, the top quarks were decayed in MadSpin [67] to W bosons that decay to $\tau$ or $\mu$ leptons, denoted commonly as $\ell$. In order to enhance the generation efficiency, we applied a generator level cut on the di-$\ell$ invariant mass of 120 GeV. The cross section of the Drell–Yan sample is $17.37$ pb and the cross section of the $t\bar{t}$ sample is $13.49$ pb. Sufficient events were generated to have an equivalent luminosity of 138 fb-1, to match that of the proton-proton collisions data taken during Run-2 (2016-18) of the LHC.
We explored a set of several BSM models as benchmarks to evaluate the performance of our approach. The details of the generation are summarized in Table 1, where the number of events is after the generator cuts are applied. Some example Feynman diagrams of the signal processes are shown in Fig. 1. For simplicity, we set the masses of the di-$\ell$ resonance in all models to 250 GeV.
In the 2HDM model [68, 69, 70] we use the heavy Higgs particle, denoted with $\Phi$, and let it decay to an $\ell$ lepton pair. In the eVLQ model [71, 72], we let the vector-like quark, $T^{\prime}$ decay to a SM top and a BSM scalar $S^{0}$. One of the $S^{0}$ then decays to an $\ell$ lepton pair while the other decays inclusively according to branching ratios similar to a heavy Higgs boson. In the NMSSM model [73, 74, 75, 76, 77] we simulate cascade processes, where the heavy resonance X decays to a Higgs boson (H) and an unknown particle Y. Y further decays to the $\ell$ leptons and the H decays invisibly, leading to missing transverse momentum. We explored non-resonant signals such as vector-like anti-leptons [78], leptoquarks [79] and heavy neutral leptons [80]. We did not achieve significant improvements with our method on these signals, as expected since these models are not well localized to the signal region, making them difficult to identify with the CATHODE approach. We therefore did not study them further.
UFO Name Extra jets Parameters #Events
SM Drell–Yan $\ell\bar{\ell}$ up to 2 N/A 3M
SM $t\bar{t}\to W^{-}W^{+}b\bar{b}\to\ell\bar{\nu}+\bar{\ell}\nu$ up to 2 N/A 2M
2HDM $t\bar{t}\Phi\to\ell\bar{\ell}$ 0 $m_{\Phi}=250$ GeV 100k
eVLQ $T^{\prime}\bar{T}^{\prime}\to tS^{0}\bar{t}S^{0}\to\ell\bar{\ell}+\text{incl}$ 0 $m_{T^{\prime}}=1000$ GeV $m_{S}=250$ GeV 100k
NMSSM $X\to YH\to\ell\bar{\ell}+\text{inv}$ 0 $m_{X}=800$ GeV $m_{H}=400$ GeV $m_{Y}=250$ GeV 100k
Table 1: List of samples used, and the details of the production. The $\ell$ symbol denotes the $\tau$ or $\mu$ leptons.
Figure 1: Feynman diagrams of the $t\bar{t}\phi$ (left), $T^{\prime}\bar{T^{\prime}}$ (middle), and $X\to YH$ (right) signal models. The $\ell$ symbol denotes the $\tau$ or $\mu$ leptons. The $S^{0}$ decays inclusively according to its branching ratios. The $H$ boson in the $X\to YH$ model decays invisibly.
@@
Figure 2 shows the reconstructed $m_{\mathrm{vis}}$ and $m_{\mu\mu}$ distributions of the di-$\tau$ and di-$\mu$ systems respectively. While the generator level mass cut is at 120 GeV, because some of the di-$\tau$ momentum is lost to the neutrinos, the reconstructed visible mass is smoothly falling starting from 100 GeV. Histograms are normalized to unit area.
Figure 2: Invariant visible mass distribution of the di-$\tau$ system (left) and the di-$\mu$ system (right)